UT Sep 12, 2019 - Restart Night #11


OSA: J. Williams
ISp: G. Bechetti

ISA: B. Rothberg, O. Kuhn (Tucson)

Mtn Manager: M. Wagner

Technical Expert: J. Hill (Tucson)

UT 01:27 - initializing all LUCIs, humidity at 89.8% dark clouds overhead

-configured LUCI-1 for N30fieldstop (already in FPU) and N375 camera - will use Br Gamma filter so background illuminates circles

L1 0151 - 2.51 sec dark

L2 0001 - 2.51 sec dark

L1 0152 - 5 sec dark (to be used for transform collection)

L2 0002 - 5 sec dark

UT 02:16 - O. Kuhn taking internal focus sequences for both MODS. Images are mods*.20190912.*.fits. Tel stationary @ 66 deg elevation (left from test preset).
  • Dual Imaging Modes: execBinoMODS dualim1.pro dualim2.pro 01-11
  • Red-Only Imaging: execBinoMODS redim1.pro redim2.pro 12-22
  • Red-Only Grating: execBinoMODS redgr1.pro redgr2.pro 23-33
  • Dual Grating: execBinoMODS dualgr1.pro dualgr2.pro 34-.. cntl-C'd as we are opening (03:57 UT)

UT 02:25 Running up the LBCs.
  • (Usual, I think?) Blue Rotator error at startup with problems with the profibus encoder reading:
    2019/09/12   02:29:25.255469         rotator system encountered a problem: problems with profibus encoder reading (8) [src/channel/channel.c:537]
   2019/09/12   02:25:57.477861   ROTATOR       GOYA no prompt return while sending > S < [src/rotator/rotator.c:2338]    2019/09/12   02:25:55.412533   ROTATOR       Encoder problems with profibus encoder reading [src/rotator/rotator.c:458]
0   2019/09/12   02:25:55.412486   ROTATOR       Profibus board 1 device 12: equipment status 33 
  • Same Blue Rotator error after turning off and back on
  • 3rd try was the charm; Blue Rotator and all LBC systems up
  • Taking 25 biases. Dome dark. John trying out RB_Science on rm507-4. Display one image pair but crashing.

UT 03:57 - enclosure opening - authorizing LUCI-1 only

-point and collimate on Stone L1, then will configure for H-band, 5sec integtration, N30 fieldstop - replicating previous setup since preset is apparently sent at H-band and we want to avoid mismatch in transform for atmospheric refraction

UT 04:20 - starting Stone L AGW-1 transform - WFS/Guider show 1.2" - borderline, but I (BR) think we're OK - no update to PCS config, using the rotator ZP as is
  • L1 0153 - 4 sec exposure for dark for subtraction
  • L1 0154 - first exposure
  • L1 0155 - Off image (might want to use 0153 instead)
  • L1 0156 - Star 17
  • L1 0157 - Star 60
  • L1 0158 - Star 124
  • L1 0159 - Star 156
  • L1 0160 - Star 166
  • L1 0161 - Star 146
  • L1 0162 - Star 132
  • L1 0163 - Star 112 - dome seeing gone, seeing <1" in WFS/Guider
  • L1 0164 - Star 167
  • L1 0165 - Star 101
  • L1 0166 - Star 52
  • L1 0167 - Star 44
  • L1 0168 - Star 16
  • L1 0169 - Star 13
  • L1 0170 - Star 87
  • L1 0171 - Star 160
  • L1 0172 - Star 105
  • L1 0173 - Star 159
  • L1 0174 - Star 126
  • L1 0175 - Star 42
  • L1 0176 - Star 110
  • L1 0177 - Star 32
  • L1 0178 - Star 55
  • L1 0179 - Star 68
  • L1 0180 - Star 77
  • L1 0181 - Star 162
  • L1 0182 - Star 173
  • L1 0183 - Star 43
  • L1 0184 - Star 54 (almost there....)
  • L1 0185 - Star 131
  • L1 0186 - Star 22
  • L1 0187 - Star 72
  • L1 0188 - Star 139
  • L1 0189 - Star 7 (stay on target...)
  • L1 0190 - Star 93
  • L1 0191 - star 0?
UT 05:20 - completed AGW-1 transform - LUCI-1 in safe config (LUCI-2 already in safe config)

UT 05:24 - J. Hill running some tests, BR sent active preset to Stone L (on-axis)

UT 05:29 - reconfiguring for LBTI

-signing off until later

UT 10:25 - taking over from LBTI - pointing and collimation near G191B2b

-putting MOS masks into storage

UT 10:27 - first time our pointing was terrible - we followd LBTI, suspect cause and effect

UT 10:28 - masks put away

UT 10:29 - going to a different pointing star - SX looks really out of focus

UT 10:42 - still working on pointing

UT 10:48 - moving to G191B2b - bino - Airmass 1.16-1.12

looks like FM4 on LUCI2 is off-center
  • L1 0192-0195 z
  • L2 0003-0006 z
  • L1 0196-0199 Ks
  • L2 0007-0010 Ks
  • L1 0200-0203 K
  • L2 0011-0014 K
  • L1 0204-0207 H
  • L2 0015-0018 H
  • L1 0208-0211 J
  • L2 0019-0022 J
UT 11:22 - moving to NS target with G200 grating zJ then HK with 1.5"/2" slit

UT 11:25 - LUCI-2 MOS successfully grabbed 1.5/2" slit and put into FPU

UT 11:42 - starting spectroscopic zJ
  • L1 0217-0220 zJ
  • L2 0028-0031 zJ
  • L1 0221-0224 HK
  • L2 0032-0035 HK
-setting GCS exposure time to 5seconds

UT 11:56 - moving to solar analog

LUCI-2 skylines look out of focus
  • L1 0228-0231 zJ
  • L2 0039-0042 zJ
  • L1 0232-0235 HK
  • L2 0043-0046 HK
UT 12:18 -collimation on SX problematic, WFS is vignetted on left side SH spots - 232-233 for HK LUCI-1 - DX does not show this problem, spots are well centered, no vignetting

-could be from sunlight comng through left vent door

UT 12:21 - clearing active optics on M1 SX

UT 12:24 - repeating HK observations
  • L1 0236-0239
  • L2 0047-0050
UT 12:31 - closing up enclosure, will take calibrations with LUCI

UT 12:34 - LUCI-1 defaulted to 1micron CWL after simply putting in blind filters

Imaging Flats - took J,H,K for both LUCIs on UT 20190906, need z and Ks

MODS cals:
  • grlamps red-only MODS1R and MODS2R: 34-36
  • grlamps blue-only MODS1B and MODS2B: 12-14
  • 5" slit flats for red-only mode: MODS1R and MODS2R: 37, 38-41 (library 10.11 has nimgs=1, I think)
  • 5" slit flats for blue-only mode: MODS1B: 15-19 and 20-22 (a comm glitch led to running "go" 2x and MODS2B: 15-17,18-20
  • red imaging flats at r, i and z: MODS1R: 44-48 (r) and MODS2R: 42-46 (r)
    • New Dust or Debris Donuts? have a look at MODS2R 0042 - are these two new donuts?
    • A dual imaging mode pair: mods2b 21 and mods2r 44 do not show these two new features.
    • Red-Only imaging: mods2r 45 (r), 46 (i) and 47 (z). The spots are on the r, i and z images. I suspect they are from something on the red folding flat on the dichroic drum, since it is the only red channel optic that the dual and red-only modes do not have in common. But the mirror is Ag coated (cleaning should await clearance from Eng & OSU).

LUCI NS observation was:
  • L1: 212-216 acquisition, 217-220 zJ, 221-224 HK
  • L2: 023-027 acquisition, 028-031 zJ, 032-035 HK
LUCI solar-analog was:
  • L1: 225-227 acquisition, 228-231 zJ, 232-235 and then 236-239 HK (repeated HK because of collimation problems).
  • L2: 036-038 acquisition, 039-042 zJ, 043-046 and then 047-050 HK
Calibrations the NS target are:
  • Flats HK
    • took darks two times as we had to repeat script because I had left 1" mask in the cal script for L1
    • L1: 240-249 are darks for continuum flats, 250-254 use Halo2
    • L2: 51-60 are darks, 61-65 use Halo2
    • L1: 255-260 are arcs
    • L2: 066-071 are arcs
  • Flats zJ
    • L1: 261-265 are darks, 266-270 use Halo2
    • L2: 72-76 are darks, 77-81 use Halo2
    • L1: 271-274 are arcs
    • L2: 82-85 are arcs

Taking Imaging Flats z,J,H,K,Ks (no issue moving MOS masks to storage)
  • L1 0275-
  • L2 0086-


OUTLINE PLAN

Plan rolled over from Night #7-8-9-10
  • LUCI
    • DONE AGW-1 Transform - HIGHEST PRIORITY
      • No need to update PCS, rotator ZP looks good after astrometric mask test
    • IF spectroscopic modes are released for LUCIs
      • DONE G200 science-like OB
      • G210 science-like OB
  • DONE LBTI Checkout - min 2 hours (poor weather) - 0.5 night in good conditions
  • Test patch to GCS/PCS to address dithering NS guiding
    • LUCI
    • MODS
  • DX M3 alignment
  • LBC
    • Standard photometric fields - requires clear skies (any seeing) - required for ZP calibrations
    • Spectro-photometric standards - requires clear skies (any seeing) - required for ZP calibrations
    • Twilight Flats all filters
    • LBC sub-framing (test OBs needed)
    • NS guiding with dithers
    • Checkout observations with Bessel filters
  • MODS
    • Test NS guiding with dithering

-- BarryRothberg - 11 Sep 2019
Topic revision: r9 - 12 Sep 2019, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback